OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -x + 4.

ganeshie8 (ganeshie8):

since product of slopes of perpendicular lines is -1, the line perpendicular to y = -x + 4 will have a slope of "1" slope = 1 point : (-2, 4) equation of line with slope "m" & passing throug (x1, y1) is : y-y1 = m(x-x1) plugin (x1, y1) and m = 1 y - 4 = 1(x+2) can you simplify further ?

OpenStudy (anonymous):

ganeshie8 i mad a mistake its y = -1/2x + 4

ganeshie8 (ganeshie8):

ok np :) can u tell slope of this line ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

ok. here is the clue : for line y = mx + c, slope = m for y = -1/2x + 4, slope = ?